OpenStudy (anonymous):

solve sqrt{x+5}+sqrt{x-3}=4

OpenStudy (anonymous):

Have you tried anything?

OpenStudy (anonymous):

\( \sqrt{x+5}+\sqrt{x-3}=4\)

OpenStudy (anonymous):

so what is the answer ?

OpenStudy (anonymous):

ahaikal that's incorrect.